a. A volume of 450 cubic centimeter is measured at a pressure of 740 mm Hg and a temperature of 25.4 oC.


What is the volume in liter? Pressure in pound per square inch and temperature in degree Kelvin?



b. What is the temperature value in degree Fahrenheit when the surrounding is experiencing a 20 degrees centigrade below zero.




1
Expert's answer
2021-07-12T03:00:05-0400

a) Given volume "=450\ cubic \ centimeter"

Pressure of "=" 740 mm Hg

Temperature"\ =" "25.4^{o}C"




Volume in litre =


We know that ,


"1\ cubic\ centimeter=0.001\ liter"


"=\dfrac{450}{1000}=0.45\ liter"



Pressure in pound per square inch "="


"1mm\ of\ mercury=0.0019\ pounds\ per\ square\ inch\"



"\dfrac{740}{14.3092}=51.715" "pounds \ per\ square\ inch"



Temperature in degree kelvin =


"25.4^{o}C=273+25.4=289.4^{o}K"



b) The formula to convert into Fahrenheit will be=


"(20^{o}\times\dfrac{9}{5})+32=68^{o}F"